# Is Our Chatbot Telling Lies? Assessing Correctness of an LLM-based Dutch Support Chatbot

### v1

@@ VI-B3 Deviations
Score ==absent$=$ = 1 55% 0% 60% 0% 12%
Score $<$ 3 64% 67% 60% 33% 21%
Score $>$ 3 36% 33% 40% 67% 79%
Score ==absent$=$ = 5 18% 0% 20% 50% 26%
TABLE V: Automated answer rating accuracy (±$\pm 1$ error margin) and percentage of what score is predicted (below/above neutral) per actual label. These conditions show how often negative-rated messages are correctly predicted as such, and positive as positive. Score 1: failure on Components Defined, Score 5: success on Verbatim Guide Defined.
VI-B4 Scenarios not working
To identify the mistakes made by the chatbot and guide future work, scores that deviate much from the ground truth are analyzed. First, a downside of using REGEX is that it can miss some components or extract unrelated text. E.g”The salary button should be clicked” Instead of expected: ”Click on: salary”. Another recurring problem is the ambiguity when detecting the heuristic Off-Context. Generally, an answer may be correct, but not for a specific exception. E.g. User asks a question about a Nurse organization, however for these organization different laws hold true. Third, relatedness and completeness influence the rating alongside truthfulness, even though the ranking is distinct for each dimension. Lastly, sometimes the LLM hallucinates correctly, when it is based on sparse information in the context. E.g. Q: ”What if I click the salary button” A:”It shows a salary overview”. Although not explicitly stated in the context, the name of the button effectively inspires this correct hallucination.
VII Threats to Validity
@@
External Validity includes threats to generalizability. We propose metrics for two message types and derive insights from AFAS. The bigger question is, who can use this work? Unlike prior work, our metrics are not linked to the financial sector and can be easily adapted to other fields. Replicating our work can be the first step in seeing feasibility. And even if our metrics do not apply to a environment, the methodology can inspire the identification of custom metrics.
VIII Discussion
The customer support team is on a challenging mission to accurately and efficiently respond to all kinds of customer queries. To support the customer support team, we started on a quest to identify the right answer to a customer query. We found that answer correctness is fundamental to a right answer and can be measured in terms of the truthfulness, relatedness, and completeness of the answer.
Learning from the common mistakes to a correct answer identified in our research, the support team now has a better understanding of how to process incoming messages and the importance of crafting questions. As a result, the support team at AFAS revisited the reformulation of the questions to get the correct answer from the chatbot. Further, while our approach is not without its flaws, AFAS has integrated our solution into their system and is testing whether some answers can be directly sent to the user. Even with detecting only 28% of the 5-star rated messages, as is currently reported in our work, we anticipate freeing up to 15,000 hours per year so that the support staff can work on more challenging queries and provide near real-time answers.
This study measured the truthfulness of a response generated by a Dutch-support chatbot using generic metrics, which can apply to companies other than AFAS. Even otherwise, our methodology of translating the human decision-making process to heuristics and then metrics can be used to identify different aspects of correctness.
A notable observation of this study is that the type of user message influences the nature of mistakes. While there were initial signs of this observation in literature [19], our process of translating how humans assess the correctness of a response to metrics makes it apparent. Since most message types follow specific patterns to construct questions, it is possible to apply a subset of heuristics (instead of all) to assess its truthfulness.
Our approach shows the viability of using custom features when the data is sparse and insufficient for training machine learning models. It allows us to remove specific mistakes using custom features. Further, our approach eliminates the need for a reference answer. This is possible since the generated answer is checked against the context provided by the help document. An advantage to this approach is that we can now assess the correctness of an answer for which a reference answer does not exist. This way, the metrics can work on the fly, and we can validate an answer before it reaches the user. However, this makes the availability of context even more important for the success and failure of the system.
Compared to most studies in English, assessing correctness in Dutch was challenging since the most widely used software packages for natural language processing are in English. For example, the Python software packages used in this work are tailored for English text. This was evident in the performance of results when we used Dutch text versus English text. There was a performance gain with the English text. We extend this recommendation to other similar explorations in regional languages to consider translating the text to English for performance gain.
@@ VIII Discussion
Our approach also has limitations in detecting mistakes. Some of the mistakes are due to the limitations of automated features, while others are hard even for the human annotator to detect. We identified three edge cases that are difficult for human annotators to detect and require a thorough knowledge of the company. These cases include situations where a superior solution exists, solutions with undesirable side effects, and the use of terms that change meaning when used in the context of AFAS. Capturing these nuances using automated features is even harder.
In the future, we propose to preprocess the text to remove irrelevant statements from the message prior to comparing it to the context. This includes general statements such as greetings.
@@
Another direction is constructing a knowledge base, learning from all incoming messages and their answers, which can supplement the information in the help documentation. This can help detect mistakes currently made as a side-effect or resulting from out-of-context. Likewise, we can develop a neural network that learns from our system to capture the subtle nuances our approach misses. Prior work on this topic has shown promising results [29, 35, 25, 26, 32, 33] and student answers [65, 67]. This way, we graduate from a workable solution to scalable solution.
IX Related Work
@@
As our research assess the correctness of content, two relevant fields are identified: Natural Language Generation (NLG) and Automated Answer Grading.
Natural Language Generation
In previous studies of NLG, human evaluation is often used [68, 34, 39, 69]. To date, several studies have investigated how the human evaluation can be automated. Focusing on quality of dialogue [31, 32, 33], or on a single messsage-answer pair. Our study can be classified under the latter, so called turn-level metrics. Much of the literature on turn-level metrics is focused on comparing embeddings of text [18, 15, 14, 27]. It either test relevancy [18, 15] or improvement [14, 27]. Improvement is measured along linguistic features like readability, syntactic style and complexity [27]. There are relatively few studies in the area of content-specific aspects. The existing research is focused on rational answers [28, 29] and helpful answers [30] rather than on correctness of answers. Roychowdhury et al. suggest a framework for a financial bot that includes confidence monitoring. This monitoring aims to identify if the LLM is hallucinating in comparison to the context, with a focus on numerical hallucinations. While we have drawn inspiration from some features to evaluate hallucinations, our approach assesses content correctness using an expanded definition of truthfulness. Further, while they propose features for safeguarding a financial bot for decision-makers, our features are focused on a metric for chatbots designed for more general content and support questions. This metric is assessed by comparing it to human evaluations. While hallucination is well researched[66] and while it overlaps with truthfulness, they are not identical. They overlap in terms of producing information that is not present in the provided document base. However, hallucinated information can be accurate, e.g., correct hallucination. And on the other hand, non-hallucinated information is not always accurate. E.g. if non-hallucinated information is present in a different context than in the help documentation. To our knowledge, correctness is mentioned once before in literature as a metric, in the research from Mehri and Eskenazi [20]. They use a LLM to rate generated answers on various dimensions, among which, correctness.
Automated Answer Grading
@@
There are a number of similarites between the field of automated metrics and automated grading in education. Where our metric grades the answer of a chatbot, these models and features predict the grade a teacher would give the answer of a student [45, 16, 70, 67, 71, 17, 46]. Kumar et al. [16] discusses numerous features used in both automatic essay grading and short answer grading. Some of these features, as well as those introduced by Roy et al. [17], are either directly or indirectly incorporated into our research. However, their features are designed for training a model that is based on multiple correct answers for a single question. This approach is not directly applicable to automated metrics without using reference answers.
X Conclusions
@@
Our results inspires how the support team queries the chatbot. Further, we anticipate a gain of up to 15,000 hours per year from adopting this system. This study offers recommendations, such as how other companies can assess the correctness of their chatbot.

### v2

@@ 6.2.3 Deviations
Score ==== 1 55% 0% 60% 0% 12%
Score << 3 64% 67% 60% 33% 21%
Score >> 3 36% 33% 40% 67% 79%
Score ==== 5 18% 0% 20% 50% 26%
Table 5: Automated answer rating accuracy ($\pm 1$ error margin) and percentage of what score is predicted (below/above neutral) per actual label. These conditions show how often negative-rated messages are correctly predicted as such, and positive as positive. Score 1: failure on Components Defined, Score 5: success on Verbatim Guide Defined
6.2.4 Scenarios not working
To identify the mistakes made by the chatbot and guide future work, scores that deviate much from the ground truth are analyzed. First, a downside of using REGEX is that it can miss some components or extract unrelated text. E.g"The salary button should be clicked" Instead of expected: "Click on: salary". Another recurring problem is the ambiguity when detecting the heuristic Off-Context. Generally, an answer may be correct, but not for a specific exception. E.g. User asks a question about a Nurse organization, however for these organization different laws hold true. Third, relatedness and completeness influence the rating alongside truthfulness, even though the ranking is distinct for each dimension. Lastly, sometimes the LLM hallucinates correctly, when it is based on sparse information in the context. E.g. Q: "What if I click the salary button" A:"It shows a salary overview". Although not explicitly stated in the context, the name of the button effectively inspires this correct hallucination.
7 Threats to Validity
@@
External Validity includes threats to generalizability. We propose metrics for two message types and derive insights from AFAS. The bigger question is, who can use this work? Unlike prior work, our metrics are not linked to the financial sector and can be easily adapted to other fields. Replicating our work can be the first step in seeing feasibility. And even if our metrics do not apply to a environment, the methodology can inspire the identification of custom metrics. While the usability of the metric has not yet been tested by others in real-world settings, we evaluated our recommendations on a manually annotated training set, reflecting how the support team perceives our recommendations. In future, we can add our recommendations to the other chatbots to see how it works in practice.
8 Lessons Learned
The customer support team is on a challenging mission to accurately and efficiently respond to all kinds of customer queries. One of the most persistent challenges we faced was defining what a right answer means. We found that all these valid responses share three requirements: truthfulness, relatedness, and completeness. This definition is applicable to any support chatbot, as it uses criteria that are universal applicable across various domains and languages. These requirements provided us with a deeper understanding of the chatbot its limitations. Utilizing the definition of correctness and its requirements allows for a meaningful evaluation, whether conducted by humans or automated systems.
8.1 Scope your metric
At the outset of our research, we aimed to develop a metric capable of identifying correct answers for all incoming user questions. However, we quickly realized that it would be hard to achieve a generic metric. We discovered that not all types of questions can be treated the same. A notable observation of this study is that the type of user message influences the nature of mistakes. Initial signs of this observation in the literature (Roychowdhury et al., 2023), which gives the impression that this observation extends to other chatbots. By identifying the different types, we were able to use targeted heuristics instead of applying the same approach to everything. By narrowing the scope of our evaluation, we move away from our original goal, but we gain practical impact. It can be advised to tailor the metrics to the question types, a general evaluation does not work.
8.2 Do not aim for perfection, but aim for impact
While our approach is not without its flaws, AFAS has integrated our solution into their system and is testing whether some answers can be directly sent to the user. Table 5 shows that, among the 5-rated messages, our system successfully identified 26% as such. For the wrongly generated answers, we were able to detect 55%. Considering the annual number of inquiries (Section 2) and the fact that roughly half of all messages are either binary or instructional in nature, we estimate that detecting even 26% of the 5-star rated messages could save approximately 15,000 hours per year. This reduction in workload would allow the support staff to focus on more complex inquiries and provide users with near real-time responses.
Beyond effeciency gains, the research led to unexpected side effects. By learning from the common mistakes of the chatbot, the support team at AFAS revisited the reformulation of questions to obtain the correct answers from the chatbot. An other side effect arose in the documentation flow, since the chatbot uses the available help documentation as context, the quality of the documentation is directly related to the quality of the chatbot. A wrong answer can be caused by bugs in the chatbot, but can also be caused by mistakes in the documentation. In that sense, the chatbot is similar to a new colleague onboarding who misses documentation or reads flawed documentation (Rastogi et al., 2015; Rodeghero et al., 2021).
This emphasizes that automation and the integration of chatbots into daily workflows have a widespread impact throughout the entire process. Analyzing these chatbots can uncover systemic weaknesses within a company, such as flaws in documentation. It illustrates how automation efforts can affect broader team practices.
Analyzing chatbot mistakes and the feedback loop, helps the support team by refining their process and question crafting. In addition, We were able to exploit the feedback given on chatbot answers to improve the quality of the documentation. By constructing a feedback loop from the feedback given on chatbot answers to the documentation team we can identity parts of the help documentation that are of low quality and are most useful to improve. The chatbot serves as a tool for identifying parts of the documentation that need attention(Aghajani et al., 2020).
Organizations should be mindful of what they are evaluating. Are they evaluating the correctness of the bot, or the documentation. In addition, by integrating error analysis into support workflows and documentation review, teams can enhance both the quality of responses and the underlying knowledge base. A thorough evaluation of your chatbot can lead to improvements that extend beyond the chatbot itself.
8.3 Do not underestimate the power of custom features
Since we are working with a flexible software product that offers many features, building a dataset is challenging. It requires input from support team experts. In addition, we were dealing with specific company knowledge and nuances, which are difficult to capture using generic tools.
To address this, we experimented with custom features. We showed that by using targeted checks and leveraging AFAS its standardized documenten style, we could effectively distinguish between correct and incorrect responses. This tailored approach allows us to validate even nuances specific to the company domain. Leveraging company documentation as ground truth, serves as a powerful validator when reference answers are not available.
This approach of custom features can be leveraged if a chatbot is operating in a specialized domain knowledge or requires company specific nuances to be taken into account. Our metrics can be adapted by learning from the documentation style of a particular company, other metrics focus on textual comparison which are independent of language or business which make them even more adaptable.
We discovered that even simple targeted heuristics and well-structured documentation can improve response validation. Organizations should treat company documentation as a powerful ground truth.
@@ 8.3 Do not underestimate the power of custom features
However, our approach also has limitations in detecting mistakes. Some of the mistakes are due to the limitations of automated features, while others are hard even for the human annotator to detect. We identified three edge cases that are difficult for human annotators to detect and require a thorough knowledge of the company. These cases include situations where a superior solution exists, solutions with undesirable side effects, and the use of terms that change meaning when used in the context of AFAS. Capturing these nuances using automated features is even harder.
@@
To address these limitations, a direction is to construct a knowledge base that learns from all incoming messages and their answers. This can supplement the information in the help documentation and help detect mistakes due to overseen side effects or from out-of-context. Likewise, we can develop a neural network that learns from our system to capture the subtle nuances our approach misses. Prior work on this topic has shown promising results (Tao et al., 2018; Lowe et al., 2017; Singh et al., 2021; Yan et al., 2016; Huang et al., 2020; Pang et al., 2020) and student answers (Xue et al., 2021; Ormerod et al., 2022). This way, we graduate from a workable solution to a scalable solution.
8.4 Translate your data
Compared to most studies in English, assessing correctness in Dutch was challenging. This was evident in the performance of results when we used Dutch text versus English text; there was a performance gain with the English text. This challenge is relevant to many other regional languages, that are less studied than large languages like English. Since the most widely used software packages for natural language processing are in English. For example, the Python software packages used in this work are tailored for English text. When working in languages with weaker NLP support, translation into English can be a practical workaround to leverage stronger tools and models. We extend this recommendation to other similar explorations in regional languages to consider translating the text to English for performance gain.
8.5 Structured Approach
In this study, we evaluated the truthfulness of responses generated by a Dutch-support chatbot using simple and adaptable metrics. Our methodology translates the human decision-making process into heuristics and then into measurable metrics, providing a structured approach to assess correctness.
Since this approach is based on how evaluators naturally judge responses, it is not specific to AFAS or limited to Dutch. Any organization can employ the same process to evaluate correctness within their own domain. The value lies not only in the specific metrics but also in the methodology itself, which involves employees in defining correctness criteria and translating these into metrics.
Other organizations can replicate this process by capturing the decision-making process of their end-users and converting it into automated metrics. This makes correctness assessment both structured and adaptable to various domains and languages.
9 Related Work
@@
As our research assess the correctness of content, two relevant fields are identified: Natural Language Generation (NLG) and Automated Answer Grading. The chatbot in question will be categorized as user-driven (Følstad et al., 2019) and support bot, specifically a Generative Question Answering (GQA) bot (Ji et al., 2023a).
9.1 Natural Language Generation
In previous studies of NLG, human evaluation is often used (Adiwardana et al., 2020; Deriu et al., 2020; Goodman et al., 2023; Serban et al., 2017). To date, several studies have investigated how human evaluation can be automated. Focusing on dialogue quality (Yeh et al., 2021; Huang et al., 2020; Pang et al., 2020), or on a single message-answer pair. Our study falls under the latter, known as turn-level metrics. Much of the literature on turn-level metrics is focused on comparing text embeddings (Merdivan et al., 2020; Zhang et al., 2020; Banerjee et al., 2023; Das and Verma, 2020). It either tests relevancy (Merdivan et al., 2020; Zhang et al., 2020) or improvement (Banerjee et al., 2023; Das and Verma, 2020). Improvement is measured along linguistic features like readability, syntactic style and complexity (Das and Verma, 2020). There are relatively few studies in the area of content-specific aspects. The existing research is focused on rational answers (Phy et al., 2020; Tao et al., 2018) and helpful answers (Gupta et al., 2022) rather than on correctness of answers. Ji et al. (2023a) mention the concept of correctness and highlight that in the field of GQA, there is an absence of standardized definitions. They also note that current methods for assessing the factual correctness of answers usually rely on human evaluation, and better automatic evaluation is needed.
(Roychowdhury et al., 2023) suggest a framework for a financial bot that includes confidence monitoring. This monitoring aims to identify whether the LLM is hallucinating in comparison to the context, with a focus on numerical hallucinations. While we have drawn inspiration from some features to evaluate hallucinations, our approach assesses content correctness using an expanded definition of truthfulness. Further, while they propose features for safeguarding a financial bot for decision makers, our features are focused on a metric for chatbots designed for more general content and support questions. This metric is assessed by comparing it to human evaluations.
While hallucination is well researched (Huang et al., 2023) and overlaps with truthfulness, they are not identical. Both involve producing information not present in the document base. However, hallucinated information can sometimes be accurate (e.g., correct or extrinsic hallucination (Ji et al., 2023a).) On the other hand, non-hallucinated information is not always accurate. E.g., if non-hallucinated information is present in a different context than in the help documentation.
Correctness is mentioned before in literature as a metric, in the research from Mehri and Eskenazi (2020). They use a LLM to rate generated answers on various dimensions, among which, correctness. Using an LLM in automated evaluation effectively assesses coherence and consistency (Ke et al., 2022). However, this method has been found to be costly and inconsistent (Lin and Chen, 2023). To address these issues, we explored the use of deterministic methods.
9.2 Automated Answer Grading
@@
There are a number of similarites between the field of automated metrics and automated grading in education. Where our metric grades the answer of a chatbot, these models and features predict the grade a teacher would give the answer of a student (Jamil and Hameed, 2023; Kumar et al., 2019; Mukti et al., 2023; Ormerod et al., 2022; Lakshmi and Simha, 2022; Roy et al., 2016; Vij et al., 2019). Kumar et al. (2019) discusses numerous features used in both automatic essay grading and short answer grading. Some of these features, as well as those introduced by Roy et al. (2016), are either directly or indirectly incorporated into our research. However, their features are designed for training a model that is based on multiple correct answers for a single question. This approach is not directly applicable to automated metrics without using reference answers.
10 Conclusions
@@
Our results inspires how the support team queries the chatbot. Furthermore, we estimate a gain of up to 15,000 hours annually through accurately identifying incorrect generated answers. Our study also offers recommendations to chatbot-building software companies. With this approach they can improve the quality of their chatbot by implementing guardrails that prevent the bot from telling lies. These faulty answers can be used as trainings data to improve the answers, or can be simply marked and not be returned to the customer. We have focused on the correctness of answers and our approach thus should be extended to other characteristics as well. Still we believe that this study serves as inspiration for comparable chatbot systems.
The proposed definition of correctness provides a structured basis for human feedback. The three dimensions truthfulness, relatedness and completeness enable deeper insights into the shortcomings of a chatbot.
When aiming for an automated metric, our methodology can be adopted, which is not restricted to a particular domain. Take the automotive industry, a mechanic and a salesman may both query a chatbot about the same car. Yet, their need of information will differ and thus their evaluation. The decision tree helps capture their way of thinking, which can then guide the design of heuristics and metrics.
The metrics we suggest can be used as a solid first step for building quick, custom measures, while the overall method gives a practical foundation for anyone who wants to design automated metrics that fit their own chatbot.
Contributions
All authors have contributed to the conception and design of the study. The study originates from the MSc Thesis of Herman Lassche. He performed material preparation, data collection, and analysis, with major contributions from Michiel Overeem and Ayushi Rastogi. Michiel Overeem was mainly involved in material preparation and data collection, while Ayushi Rastogi contributed primarily to the analysis. Both also assisted in other parts of the process. Herman Lassche wrote the initial draft of the manuscript, and all authors provided feedback during iterations on this draft. Every author has read and approved the final manuscript.
Data and Materials Availability
The data contains sensitive information about AFAS’s customers and AFAS itself. Therefore, it cannot be shared. Anonymizing the data is not possible, as it would still allow processes, questions, etc., to be traced back to the customer. This makes replication more challenging; hence, we have included a dummy data file in our Lassche et al. (2024). All code used in the study is available in our replication package. In addition, more comprehensive versions of the tables can be found in that package, highlighting more features examined as investigated in the thesis.
Usage of AI Tools
During the preparation of this work the authors used ChatGPT (mainly GPT-4o) in order to rephrase sentences and check for language mistakes. After using this tool/service, the authors reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Funding
During the research, Herman Lassche received an internship fee from AFAS. After the research, he became employed at AFAS and receives a salary. Michiel Overeem also receives a salary from AFAS. All needed materials and licenses are funded by AFAS. Ayushi Rastogi has no financial links to AFAS. No additional funding is involved.